OpenStudy (anonymous):

Solve for x: √(13+ √x) = √x+1

hartnn (hartnn):

Square both sides.

OpenStudy (anonymous):

\[t=\sqrt{x}\]and square both sides

OpenStudy (anonymous):

but i cant seem to factorise at the end...

OpenStudy (anonymous):

\[\sqrt{13+\sqrt{x}}=1+\sqrt{x}\]\[\sqrt{13+t}=1+t\]

OpenStudy (jiteshmeghwal9):